### v1

# Baryon number violation with top quark operators in the SMEFT

Abstract
Within the Standard Model Effective Field Theory framework, we set indirect constraints on top quark operators that violate baryon number by one unit above the TeV scale. We find that these constraints are typically many orders of magnitude more stringent than the recently derived direct bounds from collider experiments. Therefore, direct observation of baryon number violation in top-quark observables at the TeV scale would imply a large fine-tuning among operators across different energy scales. This possibility is not protected under universal radiative corrections or any known symmetry principles.
Introduction. The conservation of Baryon Number ($B$) is crucial for the stability of matter in the universe Weyl (1929). However, Baryon Number Violation (BNV) is indispensable for explaining the observed baryon asymmetry, a necessary ingredient for successful baryogenesis Sakharov (1967). $B$ conservation, which appears as an accidental symmetry in the Standard Model (SM), is broken by higher-dimensional operators built from the same dynamical fields and symmetry principles, suppressed by an unknown heavy scale Weinberg (1979).
Grand Unified Theories (GUTs) suggest that BNV should manifest through processes such as proton decay, accommodating extraordinarily long proton lifetimes on the order of $10^{30}$ years or more. Experimental efforts, including those at underground detectors like Super-Kamiokande, actively search for proton decay, which would provide critical evidence in favor of these unified theories Nath and Fileviez Perez (2007).
Expanding the scope of BNV studies beyond traditional searches, recent research has explored interactions involving heavier fermions. Ref. Hou et al. (2005) examines operators involving third-generation flavors. Tau-related BNV processes are discussed in Refs. Hou et al. (2005); Crivellin and Hoferichter (2023); Heeck and Watkins (2024), while BNV effects in bottom physics are explored in Ref. Beneke et al. (2024).
In the top-quark sector, a study by the CMS collaboration using proton-proton collision data at $\sqrt{s}=13$ TeV and an integrated luminosity of 138 fb-1 investigated BNV in top quark production and decay processes Hayrapetyan et al. (2024). This study provides the most stringent direct constraints to date on BNV top quark operators, probing Wilson coefficients as small as $0.7-0.02$ TeV-2 depending on the fermion flavor combination, thus improving previous bounds by three to six orders of magnitude.
The top quark stands out among SM fermions due to its heavy mass. Beyond tree level, the top-quark Yukawa coupling $y_{t}$ has motivated numerous Standard Model Effective Field Theory (SMEFT) studies that highlight top-quark contributions, especially involving interactions conserving baryon number Fox et al. (2008); Grzadkowski and Misiak (2008); Drobnak et al. (2012); Brod et al. (2015); Altmannshofer et al. (2023); Cirigliano et al. (2016); Bißmann et al. (2020); Barducci et al. (2018); Maltoni et al. (2019); Brivio et al. (2020); Durieux et al. (2019); Hartland et al. (2019); Bißmann et al. (2021); Bruggisser et al. (2021); Ethier et al. (2021); Miralles et al. (2022); Durieux et al. (2022); de Blas et al. (2022); Bruggisser et al. (2023); Giani et al. (2023); Kassabov et al. (2023); Grunwald et al. (2023); Garosi et al. (2023); Endo et al. (2019); Aebischer and Kumar (2020); Vale Silva (2024). Regarding BNV, Ref. Dong et al. (2012) not only focuses on direct LHC bounds, but also highlights the tree-level contributions from $udt\ell$ operators, with $\ell=e,\mu$, where $W$ exchange drives BNV nucleon decay. Ref. Beneito et al. (2024) discusses BNV mediated by dimension-7 operators involving light flavors, which receive a $y_{t}^{2}$ enhancement through top-loop contributions to the wave function renormalization of the scalar leg Liao and Ma (2016).
In this letter, we use the SMEFT framework to derive model-independent limits on dimension-6 top quark BNV operators from nucleon decay lifetimes. We provide a comprehensive analysis, identifying the internal mechanisms that induce low-energy BNV processes from top quark operators and find that constraints from nucleon lifetimes are typically 20 orders of magnitude more stringent than direct collider limits Hayrapetyan et al. (2024). Thus, our results suggest that directly observing BNV in top-quark processes at the TeV scale through collider experiments would generally require extreme fine-tuning among operators across different low-energy processes and energy scales.
Low-energy observables. The observables considered in our analysis are the nucleon lifetimes for decays where a nucleon ($N=p,n$) decays into a pseudoscalar ($P=\pi,\,K,\,\dots$) and a lepton ($\mathcal{L}=\ell,\,\nu_{\ell}$ with $\ell=e,\mu$), indicating $\Delta B=1$ BNV. These decays provide the most stringent and theoretically clean constraints. We summarize them in Table 1.
Channel Limit [$10^{30}$ years]
$p\to\pi^{0}e^{+}$ $2.4\times 10^{4}$ Takenaka et al. (2020)
$p\to\pi^{0}\mu^{+}$ $1.6\times 10^{4}$ Takenaka et al. (2020)
$p\to\pi^{+}\bar{\nu}$ $3.9\times 10^{2}$ Abe et al. (2014a)
$p\to K^{0}e^{+}$ $1.0\times 10^{3}$ Kobayashi et al. (2005)
$p\to K^{0}\mu^{+}$ $4.5\times 10^{3}$ Matsumoto et al. (2022)
$p\to K^{+}\bar{\nu}$ $5.9\times 10^{3}$ Abe et al. (2014b)
$n\to\pi^{-}e^{+}$ $5.3\times 10^{3}$ Abe et al. (2017)
$n\to\pi^{-}\mu^{+}$ $3.5\times 10^{3}$ Abe et al. (2017)
$n\to\pi^{0}\bar{\nu}$ $1.1\times 10^{3}$ Abe et al. (2014a)
$n\to K^{0}\bar{\nu}$ $1.3\times 10^{2}$ Kobayashi et al. (2005)
Table 1: Limits on nucleon lifetimes for various decay channels (all at 90% C.L.).
The low-energy description of these decays is given by the Low-energy Effective Field Theory (LEFT), where the heavy SM degrees of freedom have been integrated out. In the current state-of-the-art, the running and matching of dimension-6 LEFT operators are known up to one-loop order Jenkins et al. (2018a, b); Dekens and Stoffer (2019), and the branching ratios in this basis can be readily obtained using the results from Ref. Nath and Fileviez Perez (2007) (see also Ref. Beneito et al. (2024) for $p\to K^{0}\ell^{+}$).
From top operators to light-quark transitions. Given the energy scale of nucleon decay, the LEFT provides a complete description of the corresponding BNV processes, assuming they are induced by heavy particles, such as leptoquarks Dorsner et al. (2012). Top quarks are not active degrees of freedom at low energies and, consequently, there are no top-quark operators in the LEFT framework. However, the presence of top quarks, like any other high-energy particle, leaves a measurable imprint on low-energy dynamics, and operators involving top quarks are no exception.
Assuming BNV is induced by particles above the electroweak scale, these effects can be studied in a model-independent manner using the SMEFT framework, including all dimension $D>4$ operators with SM fields that preserve the SM gauge symmetries, $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$. In this context, one might consider top-quark operators such as $\varepsilon_{\alpha\beta\gamma}\,d_{R}^{\alpha}Cu_{R}^{\beta}\;t_{L}^{\gamma}Ce_{L}$, where $C$ denotes the Dirac matrix for charge conjugation and the Greek letters indicate $SU(3)_{C}$ color indices. However, this operator is not invariant under $SU(2)_{L}$ and is therefore not a valid SMEFT operator. To achieve gauge invariance, it must be extended to $\varepsilon_{\alpha\beta\gamma}\,d_{R}^{\alpha}Cu_{R}^{\beta}\;(t_{L}^{\gamma}Ce_{L}-b_{L}^{\gamma}C\nu_{L})$, naturally connecting processes with and without tops in a model-independent way. A complete basis of $\Delta B=1$ BNV dimension-6 SMEFT operators is given in Refs. Abbott and Wise (1980); Alonso et al. (2014),
$\mathcal{Q}_{prst}^{duq\ell}$ $=\varepsilon_{\alpha\beta\gamma}\,\varepsilon_{ij}\,(d_{p}^{\alpha}\,C\,u_{r}^{\beta})\,(q_{s}^{i\gamma}\,C\,\ell_{t}^{j})~{},$ (1)
$\mathcal{Q}_{prst}^{qque}$ $=\varepsilon_{\alpha\beta\gamma}\,\varepsilon_{ij}\,(q^{i\alpha}_{p}\,C\,q_{r}^{j\beta})\,(u^{\gamma}_{s}\,C\,e_{t})~{},$ (2)
$\mathcal{Q}_{prst}^{qqq\ell}$ $=\varepsilon_{\alpha\beta\gamma}\,\varepsilon_{il}\,\varepsilon_{jk}\,(q_{p}^{i\alpha}\,C\,q_{r}^{j\beta})\,(q_{s}^{k\gamma}\,C\,\ell_{t}^{l})~{},$ (3)
$\mathcal{Q}_{prst}^{duue}$ $=\varepsilon_{\alpha\beta\gamma}\,(d^{\alpha}_{p}\,C\,u_{r}^{\beta})\,(u^{\gamma}_{s}\,C\,e_{t})~{}.$ (4)
Here, $q$ and $\ell$ represent the left-handed doublets of quarks and leptons, respectively, with $q=(u_{L},d_{L})^{T}$ and $\ell=(\nu_{L},e_{L})^{T}$. The symbols $u$, $d$, and $e$ are used for the right-handed fermions corresponding to up-type quarks, down-type quarks, and charged leptons, respectively. Roman letters $i$ to $l$ refer to $SU(2)_{L}$ indices, while $p$ to $t$ denote flavor (generation) indices ranging from 1 to 3. These operators (1)-(4) form a closed set under the renormalization group equations (RGEs) Alonso et al. (2014).
The nucleon decays introduced previously are mediated by light-quark LEFT operators in the mass basis, $u^{\prime}{}_{L,R}^{1}$ and $d^{\prime}{}_{L,R}^{1,2}$. However, the quark fields in SMEFT are defined in the flavor basis, where the Yukawa matrices are not diagonal. In a general setup, these bases are related by a unitary transformation in flavor (generation) space,
$u_{L}=U_{L}^{u}u^{\prime}_{L}\,,\quad d_{L}=U_{L}^{d}d^{\prime}_{L}\,,$ (5)
and analogous ones for the right-handed fields. In our previous example,
$\mathcal{Q}_{113\ell}^{duq\ell}$ $=\varepsilon_{\alpha\beta\gamma}\,(d_{R}^{1\alpha}\,C\,u_{R}^{1\beta})(u_{L}^{3\gamma}Ce_{L,\ell}-d_{L}^{3\gamma}C\nu_{L,\ell})$
$\supset\varepsilon_{\alpha\beta\gamma}\;U_{R,11}^{u}U_{R,11}^{d}\;(d^{\prime}{}_{R}^{1\alpha}\,C\,u^{\prime}{}_{R}^{1\beta})$
$\cdot(U_{L,31}^{u}\,u^{\prime}{}_{L}^{1\gamma}Ce_{L,\ell}-U_{L,31}^{d}\,d^{\prime}{}_{L}^{1\gamma}C\nu_{L,\ell})\,.$ (6)
Due to the left-handed fields being part of $SU(2)_{L}$ doublets, it is not possible to simultaneously diagonalize both $Y_{u}$ and $Y_{d}$. Consequently, one cannot choose a flavor basis where both $u_{L}^{p}=u^{\prime}{}_{L}^{p}$ and $d_{L}^{p}=d^{\prime}{}_{L}^{p}$ hold simultaneously. Therefore, as illustrated in the example above, top-quark operators with $t_{L}\subset q^{3}$ in any flavor basis will generally induce light-quark operators in the mass basis at tree level, leading to nucleon decay.
There is, however, the freedom to select a flavor basis in which $Y_{u}$ is diagonal. In such a basis, one might focus on operators involving only $t_{R}=t^{\prime}_{R}$, to avoid the induction of light quarks at tree level. The first important caveat is the instability of the diagonality of $Y_{u}$ under SM running, as emphasized in Ref. Aebischer and Kumar (2020). If nature is selective enough to couple only to $t_{R}$ in a basis where $Y_{u}$ is diagonal, it appears more natural for this to occur at a high UV scale $\Lambda_{\text{UV}}$. However, even if that were the case, the up-type Yukawa matrix will no longer be diagonal at lower energy scales, such as the electroweak scale $\Lambda_{\text{EW}}$, due to contributions from the non-diagonal $Y_{d}$ matrix. The back rotation required to re-diagonalize $Y_{u}$ will then inevitably induce light-quark operators. Analytically, one finds for $t_{R}$ (analogously for $c_{R}$), in the leading-log approximation,
$t_{R}\to-\frac{3}{2}\frac{\epsilon_{\pi}\,y_{t}\,y_{u}}{y_{t}^{2}-y_{u}^{2}}\ln\frac{\Lambda_{\text{UV}}^{2}}{\Lambda_{\text{EW}}^{2}}\sum_{k=d,s,b}V_{tk}\,y_{k}^{2}\,V_{uk}^{*}\,u^{\prime}_{R}\,,$ (7)
where $V$ is the Cabibbo-Kobayashi-Maskawa (CKM) matrix and $\epsilon_{\pi}\equiv 1/(4\pi)^{2}$; see the supplementary material for its derivation. This effect can still be evaded if the UV model is defined in a basis where the Yukawa matrix is diagonal at $\Lambda_{\text{EW}}$. From this point onward, we will work under this assumption.
We now turn to the second important caveat. From the perspective of the flavor-to-mass basis rotation, the setup described above avoids the introduction of light quarks in the first approximation. However, pure SM interactions contain mechanisms that convert $t_{R}$ into light quarks, regardless of the choice of basis. As a result, interactions involving only light-quark generations inevitably emerge if an operator with $t_{R}$ is introduced at a high energy scale. Indeed, using the SMEFT running computed in Ref. Alonso et al. (2014), the corresponding top-philic setup is not stable under the universal running effects of BNV operators. The primary mechanism driving operator mixing involves Higgs exchanges, as illustrated in Fig. 1.
[…]
Figure 1: An example of Feynman diagram illustrating the SMEFT running of a BNV operator (black square vertex). The red vertices denote the Yukawa couplings, connected by the dashed line representing the Higgs field. The black lines correspond to light quarks, while the blue line represents a top quark, which is converted into a light quark via Yukawa insertions.
As an example, $C^{duq\ell}_{131\ell}$ appears in the $\beta$ function of $C^{qqq\ell}_{131\ell}$. Thus, even if $C^{qqq\ell}_{131\ell}(\Lambda_{\text{UV}})=0$,
$C^{qqq\ell}_{131\ell}(\Lambda_{\mathrm{EW}})\propto C^{duq\ell}_{131\ell}(\Lambda_{\mathrm{UV}})\,\epsilon_{\pi}\ln\frac{\Lambda_{\mathrm{UV}}^{2}}{\Lambda_{\mathrm{EW}}^{2}}+\cdots\,,$ (8)
BNV effects are consequently induced in tree-level nucleon decays. In this particular example, a third-generation $SU(2)_{L}$ doublet is generated at $\Lambda_{\text{EW}}$, which necessarily includes light flavors. We have verified that most operators not directly probed at tree level are tested through this mechanism. The few remaining operators produce effects via one-loop $\beta$ functions, arising from terms of the form $\left(\epsilon_{\pi}\ln\Lambda_{\mathrm{UV}}^{2}/\Lambda_{\mathrm{EW}}^{2}\right)^{2}$ rather than the single logarithm in Eq. (8).
The model-independent mechanisms described here inducing nucleon decays carry some suppression, either due to small CKM angles in the flavor-to-mass basis rotation, loop factors, or small Yukawa couplings. However, this suppression is largely overcompensated by the superior sensitivity of nucleon decay experiments compared to direct BNV searches at the LHC. As a result, indirect bounds on top-quark operators are going to be typically many orders of magnitude stronger than the direct searches.
Bounds on top BNV SMEFT operators. In this section, we present an analysis of indirect bounds on top-quark operators. As argued previously, indirect constraints are generally less stringent in the up flavor basis, where the up quark Yukawa matrix is diagonal at $\Lambda_{\text{EW}}$, with $q=(u_{L},Vd_{L})^{T}$. We work in this basis and generate the top operators at the UV scale $\Lambda_{\text{UV}}=1\,\mathrm{TeV}$. Our analysis includes not only the key effects discussed above but also the full dimension-6 SMEFT one-loop running from the TeV scale down to the electroweak scale Alonso et al. (2014), the one-loop matching to the dimension-6 LEFT Dekens and Stoffer (2019), and the LEFT running Jenkins et al. (2018a) down to the low-energy scale. In practice, we use the Mathematica package DsixTools Celis et al. (2017); Fuentes-Martin et al. (2021). Additional technical details on our specific setup are provided in the supplementary material.
Once the branching ratios are obtained as functions of the SMEFT Wilson coefficients at the TeV scale, we fit them to the current experimental limits in Table 1, including one Wilson coefficient at a time. In the few cases where redundancies arise, such as $\mathcal{Q}^{qque}_{1311}=\mathcal{Q}^{qque}_{3111}$, we adopt the convention of inducing the corresponding Wilson coefficients with the same strength. For instance, in this example, we display only the upper experimental bound on $C^{qque}_{1311}$, assuming that in the original $1\,\mathrm{TeV}$ Lagrangian, all Wilson coefficients are zero except for $C^{qque}_{3111}=C^{qque}_{1311}$. Once again we relegate further details to the supplementary material.
The experimental limits on all Wilson coefficients associated with top operators are displayed in Fig. 2. In all cases, these bounds are several orders of magnitude more stringent than the corresponding direct limits from Ref. Hayrapetyan et al. (2024).
Figure 2: Constraints on $\Delta B=1$ SMEFT Wilson coefficients (1)-(4) (red, purple, blue, and green, respectively) at 68% C.L., in TeV-2 units, derived from limits on $N\to P\,\bar{\mathcal{L}}$ nucleon lifetimes, see Table 1.
To better understand the hierarchy of the obtained bounds, we repeat the analysis, retaining only the tree-level SMEFT-LEFT matching factors and the leading-log effects. At this stage, a comment is in order. In a new physics scenario where UV couplings are of order one, with $\Lambda_{\mathrm{BNV}}\sim 10^{13}\,\mathrm{TeV}$, truncating the leading logarithmic series (scaling as powers of $\epsilon_{\pi}\ln\Lambda_{\mathrm{UV}}^{2}/\Lambda_{\mathrm{EW}}^{2}$ multiplied by order one couplings such as the gauge and top Yukawas) may be unjustified. In such cases, one would need either to perform the full one-loop RGE running, as we have implemented numerically above, or to solve it analytically, resumming at least the problematic logarithmic series with order one couplings. However, within our setup, where top operators are generated at the TeV scale, the logarithms are large but truncating the series remains a safe approximation.
Within the setup described above, we determine the parametric suppression of the prefactor that numerically dominates the experimental bound for each Wilson coefficient, which for simplicity we assume to be real. The corresponding results are presented in Table 2. Notice that $V_{31}(Y_{d,31})$ and $V_{13}(Y_{d,31})$ include a non-negligible CP-violating phase, which we take into account.
$\boldsymbol{C^{duq\ell}}$ $\boldsymbol{C^{qque}}$
1 1 $1\,1\,3\,\ell$ $V_{32}$ $p\to K^{+}\bar{\nu}$ 1 3 1 113111\,3\,1\,11 3 1 1 $V_{31}$ $p\to\pi^{0}e^{+}$
2 1 $2\,1\,3\,\ell$ $V_{31}$ $p\to K^{+}\bar{\nu}$ 1 3 1 213121\,3\,1\,21 3 1 2 $V_{31}$ $p\to\pi^{0}\mu^{+}$
[…]
$\boldsymbol{C^{qqq\ell}}$ $\boldsymbol{C^{duue}}$
[…]
1 3 $1\,3\,1\,\ell$ $V_{11}\,V_{32}$ $p\to K^{+}\bar{\nu}$ 1 1 $1\,1\,3\,a$ $(Y_{e})_{aa}\,(Y_{u})_{33}\,V_{32}\,L$ $p\to K^{+}\bar{\nu}$
2 3 $2\,3\,1\,\ell$ $V_{22}\,V_{31}$ $p\to K^{+}\bar{\nu}$ 1 3 1 113111\,3\,1\,11 3 1 1 $(Y_{d})_{11}\,(Y_{u})_{33}\,V_{31}\,L$ $p\to\pi^{0}e^{+}$
1 2 $1\,2\,3\,\ell$ $V_{21}\,V_{32}$ $p\to K^{+}\bar{\nu}$ 1 3 $1\,3\,1\,a$ $(Y_{e})_{aa}\,(Y_{u})_{33}\,V_{32}\,L$ $p\to K^{+}\bar{\nu}$
1 3 $1\,3\,2\,\ell$ $V_{31}\,V_{22}$ $p\to K^{+}\bar{\nu}$ 2 1 3 121312\,1\,3\,12 1 3 1 $(Y_{d})_{12}\,(Y_{u})_{33}\,V_{31}\,L$ $p\to\pi^{0}e^{+}$
1 3 $1\,3\,3\,\ell$ $g^{2}\,V_{31}\,V_{32}\,L$ $p\to K^{+}\bar{\nu}$ 2 1 $2\,1\,3\,a$ $(Y_{e})_{aa}\,(Y_{u})_{33}\,V_{31}\,L$ $p\to K^{+}\bar{\nu}$
[…]
2 2 $2\,2\,3\,\ell$ $(Y_{d})_{23}\,(Y_{d})_{13}\,V_{21}\,V_{32}\,L$ $p\to K^{+}\bar{\nu}$ 2 3 1 323132\,3\,1\,32 3 1 3 $(Y_{e})_{33}\,(Y_{u})_{33}\,V_{31}\,L$ $p\to K^{+}\bar{\nu}$
3 2 $3\,2\,3\,\ell$ $(Y_{d})_{33}\,(Y_{d})_{13}\,V_{21}\,V_{32}\,L$ $p\to K^{+}\bar{\nu}$ 3 1 $3\,1\,3\,c$ $(Y_{d})_{13}\,(Y_{u})_{33}\,V_{31}\,L$ $p\to\pi^{0}\ell^{+}$
2 3 $2\,3\,2\,\ell$ $(Y_{d})_{33}\,(Y_{d})_{13}\,V_{21}\,V_{22}\,L$ $p\to K^{+}\bar{\nu}$ 3 3 $3\,3\,1\,c$ $(Y_{d})_{13}\,(Y_{u})_{33}\,V_{31}\,L$ $p\to\pi^{0}\ell^{+}$
Table 2: Parametric suppression of leading effects constraining top-quark BNV operators for different Wilson coefficients at $\Lambda_{\mathrm{UV}}$. The table is divided into four blocks, each corresponding to a specific Wilson coefficient: $C^{duq\ell}$, $C^{qqq\ell}$, $C^{qque}$, and $C^{duue}$, respectively. For each set, the relevant indices ($c=1,2$ ; $a=2,3$ ; $i,\ell=1,2,3$) and parametric suppression ($L\equiv\epsilon_{\pi}\ln\Lambda_{\mathrm{UV}}^{2}/\Lambda_{\mathrm{EW}}^{2}$) are displayed in the first and second columns; $g$ corresponds to the electroweak coupling of the SM. The most constraining nucleon decay channels are indicated in the last column. ${}^{*}\,$ote that $C^{duq\ell}_{3132}$ is dominated by the $(Y_{d})_{13}\,(Y_{e})_{22}\,V_{31}\,L$ prefactor entering the experimental $p\to\pi^{0}\mu^{+}$ channel.
Using these tables, the overall hierarchy in Fig. 2 becomes clearer. Wilson coefficients of top operators that contribute to nucleon decays at tree level, suppressed only by CKM angles, are the most tightly constrained. Bounds on operators appearing at the leading-log level and involving light Yukawa insertions are relatively less stringent. When lepton Yukawa insertions are involved, operators with heavy leptons are more strongly constrained. Finally, the few Wilson coefficients that do not appear at order $\epsilon_{\pi}\ln\Lambda_{\mathrm{UV}}^{2}/\Lambda_{\mathrm{EW}}^{2}$ are the least constrained within the mechanisms studied here, since they only emerge at order $\left(\epsilon_{\pi}\ln\Lambda_{\mathrm{UV}}^{2}/\Lambda_{\mathrm{EW}}^{2}\right)^{2}$. In this sense, the tables also help identify cases where generally subleading effects, not considered here, could become relevant or even dominant. For instance, this applies to unknown loop contributions in the matching involving higher-dimensional effects in the LEFT framework (sometimes only suppressed by additional light Yukawa insertions) or unaccounted higher-order threshold effects in the decoupling of different quark flavors. These contributions are not enhanced by large $\ln\Lambda_{\mathrm{UV}}^{2}/\Lambda_{\mathrm{EW}}^{2}$ factors and therefore exhibit a different functional dependence compared to the effects considered here and thus no strong cancellations can be expected. Given the relevance of the decay modes $p\to K^{+}\bar{\nu}$ and $p\to\pi^{0}\ell^{+}$ in our study, a comment regarding future projections is in order. The bounds on their branching ratios are expected to improve by a factor $2$ to $5$ depending on the experiment Abe et al. (2018); Askins et al. (2020); Abi et al. (2020); Abusleme et al. (2023) (see also Ref. Domingo et al. (2024)). This improvement will enhance the sensitivity shown in Fig. 2 by a factor $\sim 2$.
Discussion and conclusions. Violation of Baryon Number has not yet been observed. Direct searches for BNV involving top operators at colliders, such as those studied in Ref. Hayrapetyan et al. (2024), have achieved remarkable precision and can be regarded as rigorous tests of SM symmetries. However, a different question is whether a nonzero result is expected in a general BSM scenario, considering the stringent low-energy bounds from nucleon decay lifetimes. The purpose of this letter has been to explore this question in a model-independent manner.
Lagrangians with operators of dimension greater than four naturally arise in generic SM extensions once the heavy degrees of freedom are integrated out. These extensions are generally described by the SMEFT framework, which includes operators involving top quarks. On the other hand, nucleon decays are low-energy processes governed by the LEFT, which does not include top-quark operators. However, the presence of top-quark operators in the SMEFT Lagrangian induces LEFT operators with some degree of suppression.
We find that, in general, the flavor basis in which the up Yukawa matrix is diagonal at the electroweak scale leads to less stringent indirect bounds on top operators. Within this setup, we conducted a comprehensive analysis using the current state-of-the-art methods to evaluate the low-energy impact of these operators, originated from new-physics dynamics at the TeV scale. Our findings show that the flavor-to-mass basis rotation, combined with one-loop SMEFT running effects (which are universal and thus independent of the UV dynamics), is sufficient to set experimental bounds on all top-quark operators. These bounds are many orders of magnitude more constraining than those obtained from direct searches.
Our results should not be interpreted as a definitive no-go theorem, as we have not conducted (and cannot conduct, given current experimental and especially theoretical limitations) a global fit that accounts for all possible operators and effects. In this context, redundancy plays an important role, and precision tests of fundamental SM symmetries, such as those in Ref. Hayrapetyan et al. (2024), are always valuable.
However, the level of fine-tuning required to avoid the nucleon decay bounds studied here is extremely difficult to conceive, particularly given the absence of stringent symmetries preventing heavy-to-light quark conversion. For instance, one might attempt to induce a light-quark operator at the TeV scale that precisely cancels the effect of the top-quark operator in the low-energy observable. Achieving this would typically require a cancellation at the level of $10-20$ decimal points among the separate contributions of different Wilson coefficients, involving different infrared logarithms between the arbitrary TeV scale and the effective one corresponding to the low-energy process. Even if such a cancellation could be arranged at tree or one-loop level, it would be completely disrupted when incorporating the next order of corrections. In this context, the fine-tuning required for such a cancellation is, in practice, impossible to calculate and, even if it were to occur for a specific low-energy BNV observable, it would generally not hold across different processes.
Thus, our work has outlined the challenges in conceiving any BSM scenario with nonzero top-quark operators at the TeV scale that could be observed in collider searches. We have done so in a transparent manner, aiming to inspire potential new approaches to overcome these challenges. For instance, one promising avenue is the exploration of BNV operators in SM extensions involving long-lived particles which cannot be produced in nucleon decays due to kinematic constraints Fornal and Grinstein (2018); Fox et al. (2024); Li et al. (2024). However, such exploration lies beyond the scope of this letter.
Acknowledgments. We thank Svjetlana Fajfer and Antonio Pich for valuable comments on the manuscript. This work is supported by the European Union – Next Generation EU, by the University of Padua under the 2021 STARS Grants@Unipd programme (Acronym and title of the project: CPV-Axion – Discovering the CP-violating axion), by MIUR grant PRIN 2017L5W2PT, by MCIN/AEI/10.13039/501100011033 Grant No. PID2020-114473GB-I00, and by PROMETEO/2021/071 (GV).